# Short-time simulation of quantum dynamics by Pauli measurements

### v1

@@ I Introduction
The study of the dynamics of quantum systems is a highly active field of research and predicting dynamical properties constitutes one of the most anticipated tasks to be tackled by quantum devices. Recent developments have seen an impressive rise in experimentally available qubit numbers and quantum operations in actual physical systems, be it for universal quantum devices [1] or sub-universal quantum simulators [2]. However, full-fledged quantum simulation is still prohibitively expensive, and purely classical methods are limited since a universal, efficient classical algorithm for quantum simulation would imply that all quantum algorithms could be efficiently simulated classically (${\rm BPP}={\rm BQP}$). This is because quantum dynamics generated by local Hamiltonians is ${\rm BQP}$ complete and hence as computationally powerful as a full quantum computer when formalized as a decision problem [3, 4, 5]. To make the best use of the currently available resources, recent developments have thus focused on finding resource-efficient algorithms, be it through hybrid quantum-classical approaches outsourcing parts of the computation to classical hardware, designing algorithms tailored to specific hardware and simulation tasks, or improving classical methods. In this delicate and important regime, trade-offs between the required circuit depth and the number of required circuit evaluations are also of key interest, since the former is highly limited in current, noisy quantum devices.
@@
Following a second line of thought, in recent years, the notion of randomized measurements has gained traction in a variety of specific flavors, following a “measure now, compute functions of interest later” mindset [6, 7, 8, 9] that is essential to the efficient functioning of most hybrid quantum-classical algorithms that can be executed on near-term hardware. Besides estimating the energy of some trial states necessary for the variational quantum eigensolver and most other variational quantum algorithms [10, 11, 12], such methods can also be used to estimate expectation values of observables with small Frobenius norms, such as overlaps with pure states [6, 13] or estimating energy gaps with no additional resources beyond time evolution and measurements [14]. These applications thus impressively showcase the fundamental power of quantum measurements and classical post-processing and highlight the need for dedicated readout schemes such as grouping strategies [15], graph optimization techniques [16] or fully randomized or de-randomized schemes [6, 7, 17, 18]. After all, the power of quantum measurement goes hand in hand with efficient measurement protocols.
@@
Randomness has found even more applications in reducing resource costs of near-term and early fault-tolerant quantum devices as it can help reduce errors that arise for product formulas [19, 20, 21, 22, 23]. Furthermore, in the context of algorithms that use what are called linear combinations of unitaries (LCU) [24, 25, 26, 27, 28, 23], randomness can help circumvent the need for block encodings [23, 29, 30]. For certain settings, these LCUs can be constructed purely in post-processing [31] without actually requiring any auxiliary qubits.
@@
While these methods approach the task of making quantum simulation more feasible from the side of improving quantum algorithms or trading resources to accommodate noisy and early fault-tolerant quantum devices, there has also been a lot of recent progress toward improved purely classical methods simulating quantum dynamics, using tensor network methods [32, 33, 34], truncated Pauli paths [35] or cluster expansions [36]. Cluster expansion methods are well-suited for the classical simulation of short-time quantum dynamics for product states evolving under a few-body Hamiltonian. They are reminiscent of the quantum algorithm simulating Hamiltonian dynamics with truncated Taylor series [27], another LCU approach, and use nested commutators to reduce computational costs. These conceptual similarities raise the crucial question of whether such methods can be used in a hybrid quantum-classical setting to help enhance current quantum devices.
@@
This work demonstrates how the power of randomized Pauli measurements can be leveraged to estimate short-time dynamics of arbitrary quantum states in classical post-processing. Moreover, this framework readily extends beyond quantum dynamics and applies to any low-order Taylor approximation of spectral functions of linear combinations of measurable operators, not just Pauli operators. As such, we anticipate further practical applications in notions of Hamiltonian learning that are based on short-time dynamics or in device verification, in short-time imaginary time evolution, or for introducing operations intractable to a given sub-universal quantum simulator in classical post-processing. These applications will be sketched here.
@@ II.1 Problem setting
The precise problem we address in this work is the following: Consider an $n$-qubit system whose Hamiltonian $H$ be specified as the linear combination of $L$ Pauli operators $H_{\ell}$ acting on an $n$-qubit system, that is,
@@
where the number of non-trivial tensor factors per Pauli operator, also denoted as their Pauli weight, is upper bounded by $w$. Furthermore, let $O$ be some (linear) observable. Then, the goal is to estimate the time dynamics $\mathrm{tr}\!\left(\mathrm{e}^{\mathrm{i}Ht}O\mathrm{e}^{-\mathrm{i}Ht}\rho\right)$ for an arbitrary, potentially mixed, initial state $\rho$ and (short-)time $t>0$.
@@ II.2 Expanding time dynamics with a Taylor series
which itself is the sum of $m\leq\sum_{k=0}^{K}L^{k}=\left(L^{K+1}-1\right)/(L-1)=\mathcal{O}\left(L^{K}\right)$ $n$-qubit Pauli operators $P_{j}$ corresponding to some $H_{\ell_{1}}\cdots H_{\ell_{k}}$ (whose weight is upper bounded by $\min\{wK,n\}$) and associated pre-factors $\beta_{j}$. As long as $\lVert H\rVert t\leq\lambda t\leq 1$, e.g., in the case of short-times, the accuracy of this approximation is upper-bounded by $\epsilon$ when choosing $K=O(\log(1/\epsilon))$. Appendix A thoroughly discusses and analyzes the truncation error of a Taylor approximation of the time evolution operator, recovering previous results of Ref. [27]. However, this is not the case anymore for longer simulation times. One way to circumvent this issue is to slice the time evolution into $r\geq\lVert H\rVert t$ segments $U(t)=U(t/r)^{r}$, each with accuracy $\epsilon/r$, increasing the number of terms in Eq. (3) by a factor of $r$ as one would for product formula approaches.
@@
which is now a sum of at most $m^{2}$ different $n$-qubit Pauli operators of the form $Q_{i}=P_{j}OP_{j^{\prime}}H_{\ell_{1}}\cdots H_{\ell_{k}}OH_{\ell_{1}}\cdots H_{\ell_{k^{\prime}}}$. A priori, the coefficients $\gamma_{i}=\beta_{j}\beta_{j^{\prime}}$ can be complex-valued, but the underlying Hermitian structure of the time-evolved observable (i.e. symmetry under taking adjoints) ensures that complex-valued coefficients must cancel out and the total sum ranges over $m_{\mathrm{tot}}\leq m^{2}$ terms with real-valued coefficients $\gamma_{i}\in\mathbb{R}$.
Estimating $\mathrm{tr}\!\left(O(t)\rho\right)$ is thus well approximated by a real-valued, linear combination of the expectation values of Pauli operators $Q_{i}$, i.e.,
@@
For $K=O(\log(r/\epsilon))$ and $r=\lVert H\rVert t$, the number of Pauli expectation values to be estimated is upper bounded by
$m_{\mathrm{tot}}\leq\left(\sum_{k=0}^{K}L^{k}\right)^{2r}\!\!=\mathcal{O}(L^{2Kr})=\mathcal{O}\left(\left(\frac{\lVert H\rVert tL}{\epsilon}\right)^{2\lVert H\rVert t}\right),$ (7)
where $\|H\|$ denotes the operator (or spectral) norm of the Hamiltonian $H$. The maximum Pauli weight over all $Q_{i}$s is given by
$w_{\mathrm{max}}=\min(\mathcal{O}\left(2ww_{O}\lVert H\rVert t\log(\lVert H\rVert t)),n\right).$ (8)
Estimating time dynamics can, therefore, effectively be reduced to a pure measurement task – relevant in many meaningful physical platforms – albeit with potentially many required measurements. Since the number of expectation values of Pauli operators to be estimated grows exponentially in the simulation time, this method is practically limited to short simulation times.
@@
It is important to note that since the Taylor expansions rely on powers of the Hamiltonian and thus powers of sums of Pauli operators, not all will be pairwise different. This is due to their group structure and the fact that Pauli operators square to the identity. Consequently, the effective number of terms in this sum can be significantly smaller than the provided upper bound. However, an exponential scaling in the chosen cutoff $K$ and number of segments $r$ cannot be fully avoided.
For time-evolved observables $O(t)$ in particular, using the fact that the time evolution in the Heisenberg picture for time-independent Hamiltonians and observables results in well-behaved nested commutators, we find that Eq. (25) can be rewritten as
@@
which is reminiscent of cluster expansion techniques [36] using nested commutators to decide which terms to estimate in a classical algorithm on product states, and stressing that we are leveraging the group structure of the Pauli operators. The advantage of such a reformulation is that we no longer have to calculate all combinations of Pauli operators to obtain the $Q_{i}$ in Eq. (5), many of which will cancel, but can use the underlying structure to calculate only non-vanishing terms. Using the knowledge that $\mathrm{tr}\!\left(O(t)\rho\right)$ has to be real-valued, we can further eliminate all odd powers of $t$. The connection to well-understood cluster expansion techniques provides a straightforward path to enhancing such techniques with quantum resources. Instead of performing such an expansion and estimating the individual Pauli expectation values classically for product states as in Ref. [36], we can measure them on arbitrary quantum states using existing expansions. Since nested commutators are a reformulation of the Taylor series of $O(t)$, the same error bounds as in Appendix B apply. Central to cluster expansion techniques and extensions thereof, thorough discussions can be found, e.g., in Refs. [36, 38]. Knowledge of the structure of the function of interest can thus help reduce the amount of classical post-processing.
Instead of concatenating two Taylor approximations of $U(t)$ for $t>0$ into an approximation of $O(t)$, as we have presented above, we can also directly Taylor expand the target function, as one would do for more generic functions $F(H,t)$. How such a direct expansion would present itself and differences in accuracy and the number of terms are discussed in Appendix B.
@@ II.3 Error bounds
An approximation of $\mathrm{tr}\!\left(F(H,t)\rho\right)$ via a truncated Taylor series followed by an estimation of the resulting linear combination of expectation values of Pauli operators, which is accompanied by two types of errors: A systematic error due to the truncation of the Taylor series and a sampling error obtained during the estimation of the Pauli expectation values. Both of these error types can be rigorously bounded.
@@ II.3.1 Systematic errors
We can upper-bound the operator norm difference between the truncated Taylor series and target matrix function $F(H,t)$ by employing Taylor’s remainder theorem as showcased in Appendix A for a direct Taylor expansion of the particular matrix-valued function $F(H,t)=O(t)$. When stitching together multiple approximations as in Eq. (4) above, where we have constructed an approximation of $O(t)$ using truncated Taylor series for $U(t)$, we can also combine the individual approximation errors to bound the total error when approximating the target function. As in Ref. [23, Lemma 1], given an approximation $\widetilde{U}$ of $U$ with $\lVert U-\widetilde{U}\rVert\leq\epsilon$, we can redefine $\widetilde{U}=U+U_{\epsilon}$ with $\lVert U_{\epsilon}\rVert\leq\epsilon$. We then find
@@ II.3.2 Sampling error
Using state-of-the-art methods, the linear combination of unitaries of Eq. (5) could be implemented deterministically using block encodings in the spirit of Refs. [37, 27] or in a randomized fashion using the frameworks of Refs. [30, 23]. However, since we are dealing with sums of Pauli operators, simply performing Pauli basis measurements as outlined in Eq. (6) suffices and no further quantum circuitry is required.
@@
where we further normalize the pre-factors $\gamma_{j}$ such that we arrive at a probability distribution $\{p_{i},P_{i}\}$. Sampling which Pauli expectation value $\mathrm{tr}\!\left(P_{i}\rho\right)$ to measure in each circuit evaluation from this distribution results in an unbiased estimator of $\widetilde{O}(t)$. Using Hoeffding’s inequality, we then find that
@@
samples suffice to estimate the linear combination of Pauli operators up to an error of $\epsilon$ with probability at least $1-\delta$.
The estimation can be improved using structural knowledge of the target function, e.g., by incorporating the nested commutator structure and quantum resources saved by immediately returning zero for all complex-valued $\gamma_{j}$ when considering $O(t)$. Alternatively, grouping strategies [15], graph optimization techniques [16], fully randomized measurements [6, 7] or other dedicated readout schemes could be used with their corresponding sampling error bounds.
@@
Especially quantum algorithms using randomized classical shadows [6, 7, 8, 9] are of additional interest. Since these strategies by default aim to estimate a set of $m$ $n$-qubit Pauli operators $P_{i}$, they constitute the prototypical “measure now, compute functions of interest later” mindset. Thus, even if the main objective of a quantum algorithm differs from those discussed here, one still has the measurement data to reuse and can formulate a sampling error bound given the number of samples.
The performances of these methods are shown mostly heuristically with an analytical upper bound on the number of measurements given by Ref. [40, Theorem 3]. It essentially states that for an accuracy $\epsilon\in(0,1)$ and failure probability $\delta$, it suffices to (sequentially) use
$N=\frac{2}{\epsilon^{2}(1-\epsilon)}3^{w}\log\frac{3L}{\delta}$ (13)
copies of $\rho$ to approximate each $\mathrm{tr}\!\left(P_{i}\rho\right)$ up to (additive) accuracy $\epsilon$, where $w$ denotes the maximum Pauli weight.
For the presented algorithm these results lead to the following conclusion: controlling the error of each $Q_{i}$ in Eq. (6) to $\epsilon/\lVert\gamma_{i}\rVert_{1}$ and thus bounding the overall sampling error of $\widetilde{O}(t)$ by $\epsilon$ requires
@@
copies of $\rho$. While the actual performance could be better by orders of magnitude, these methods usually perform best for many low-weight or few global terms in contrast to the Taylor series which outputs many terms with intermediate to large Pauli weight. Having obtained measurement data for another, primary, task, we can reuse the measurement data to secondarily estimate matrix functions $(H,t)\mapsto F(H,t)$ up to an explicitly determinable error.

### v2

@@ I Introduction
Predicting the dynamic properties of an interesting quantum system constitutes one of the most anticipated tasks to be tackled by nascent quantum computers and simulators. Recent years have seen unprecedented advances in the size (qubit number) and accuracy (gate fidelities) of such quantum hardware platforms, be it for universal quantum devices ( e.g. [1, 2, 3, 4]) or sub-universal quantum simulators ( e.g. [5, 6]). However, full-fledged quantum simulation is still prohibitively expensive and thus remains elusive. Furthermore, purely classical methods are limited since a universal, efficient classical algorithm for quantum simulation would imply that all quantum algorithms could be efficiently simulated classically (${\rm BPP}={\rm BQP}$). This is because quantum dynamics generated by local Hamiltonians is ${\rm BQP}$ complete and hence as computationally powerful as a full quantum computer when formalized as a decision problem [7, 8, 9]. Hence, a dedicated focus on resource efficiency is crucial to making the most out of currently available quantum platforms, be it through hybrid quantum-classical approaches outsourcing parts of the computation to classical hardware, designing algorithms tailored to specific hardware and simulation tasks, or improving classical methods. To this end, trade-offs between the required circuit depth and the number of required circuit evaluations are of key interest since the former is highly limited in current, noisy quantum devices.
@@
Following a second line of thought, in recent years, the notion of randomized measurements has gained traction in a variety of specific flavors, following a “measure now, compute functions of interest later” mindset [10, 11, 12, 13] that is essential to the efficient functioning of most hybrid quantum-classical algorithms that can be executed on near-term hardware. Besides estimating the energy of some trial states necessary for the variational quantum eigensolver and most other variational quantum algorithms [14, 15, 16], such methods can also be used to estimate expectation values of observables with small Frobenius norms, such as overlaps with pure states [10, 17] or estimating energy gaps with no additional resources beyond time evolution and measurements [18]. These applications thus impressively showcase the fundamental power of quantum measurements and classical post-processing and highlight the need for dedicated readout schemes such as grouping strategies [19], graph optimization techniques [20], or fully randomized or de-randomized schemes [10, 11, 21, 22]. After all, the power of quantum measurement goes hand in hand with efficient measurement protocols.
@@
Randomness has found even more applications in reducing resource costs of near-term and early fault-tolerant quantum devices as it can help reduce errors that arise for product formulas [23, 24, 25, 26, 27, 28]. Furthermore, in the context of algorithms that use what are called linear combinations of unitaries (LCU) [29, 30, 31, 32, 33, 27], randomness can help circumvent the need for block encodings [27, 34, 35]. For certain settings, these LCUs can be constructed purely in post-processing [36] without actually requiring any auxiliary qubits.
Figure 1: Cartoon illustration of our framework: We use a truncated Taylor series approach in the Heisenberg picture to achieve short-time quantum simulation entirely via classical post-processing. This trade-off between circuit depth (left) and sample complexity (right) only requires the state preparation unitary $V$ (potentially classically intractable Schrödinger evolution), as well as one parallel layer of ‘reprogrammable’ single-qubit Clifford gates (think: local Clifford/Pauli shadows).
@@
While these methods approach the task of making quantum simulation more feasible from the side of improving quantum algorithms or trading resources to accommodate noisy and early fault-tolerant quantum devices, there has also been a lot of recent progress toward improved purely classical methods simulating quantum dynamics, using tensor network methods [37, 38, 39], truncated Pauli paths [40] or cluster expansions [41]. Cluster expansion methods are well-suited for the classical simulation of short-time quantum dynamics for product states evolving under a few-body Hamiltonian. They are reminiscent of the quantum algorithm simulating Hamiltonian dynamics with truncated Taylor series [32], another LCU approach, and use nested commutators to reduce computational costs. Using backpropagated Heisenberg evolution has also sparked recent interest in the context of Pauli path methods for classically simulating quantum systems [42, 43, 44, 45]. These conceptual similarities raise the crucial question of whether such methods can be used in a hybrid quantum-classical setting to help enhance current quantum devices.
This work demonstrates how the power of randomized Pauli measurements can be leveraged to estimate short-time dynamics of arbitrary quantum states in classical post-processing, thereby trading circuit depth for additional measurements. It is inspired by viewing the results for simulating Hamiltonian dynamics with a truncated Taylor series [32] within the Heisenberg picture. The proposed method is in line with other recent insights on the power of measurements for the manipulation of quantum states [46, 47] and strengthens the “measure now, compute functions of interest later” mindset [10, 11, 12, 13].
@@
Moreover, this framework readily extends beyond quantum dynamics and applies to any low-order Taylor approximation of spectral functions of linear combinations of measurable operators, not just Pauli operators. As such, we anticipate further practical applications in notions of Hamiltonian learning that are based on short-time dynamics or in device verification, in short-time imaginary time evolution, or for introducing operations intractable to a given sub-universal quantum simulator in classical post-processing. These applications will be sketched here.
We refer to Figure 1 for a cartoon illustration of this framework and its versatility.
@@ II.1 Problem setting
The precise problem we address in this work is the following: Consider an $n$-qubit system whose Hamiltonian $H$ is specified as the linear combination of $L$ Pauli operators $H_{\ell}$ acting on an $n$-qubit system, that is,
@@
where the number of non-trivial tensor factors per Pauli operator, also denoted as their Pauli weight, is upper bounded by $w$. Furthermore, let $O$ be some (linear) observable. Then, the goal is to estimate the time dynamics $\mathrm{tr}\!\left(\mathrm{e}^{\mathrm{i}Ht}O\mathrm{e}^{-\mathrm{i}Ht}\rho\right)$ for an arbitrary (potentially mixed) initial state $\rho$ and (short) evolution time $t>0$.
@@ II.2 Expanding time dynamics with a Taylor series
which itself is the sum of $m\leq\sum_{k=0}^{K}L^{k}=\left(L^{K+1}-1\right)/(L-1)=\mathcal{O}\left(L^{K}\right)$ $n$-qubit Pauli operators $P_{j}$ corresponding to some $H_{\ell_{1}}\cdots H_{\ell_{k}}$ (whose weight is upper bounded by $\min\{wK,n\}$) and associated pre-factors $\beta_{j}$. As long as $\lVert H\rVert t\leq\lambda t\leq 1$, e.g., in the case of short times, the error of this approximation is upper-bounded by $\epsilon$ when choosing $K=O(\log(1/\epsilon))$. Appendix A thoroughly discusses and analyzes the truncation error of a Taylor approximation of the time evolution operator, recovering previous results of Ref. [32]. However, this is not the case anymore for longer simulation times. One way to circumvent this issue is to slice the time evolution into $r\geq\lVert H\rVert t$ segments $U(t)=U(t/r)^{r}$, each with error $\epsilon/r$, increasing the number of terms for the total time evolution in Eq. (3) exponentially with respect to $r$ in worst-case.
@@
which is now a sum of at most $m^{2}$ different $n$-qubit Pauli operators of the form $Q_{i}=P_{j}OP_{j^{\prime}}=H_{\ell_{1}}\cdots H_{\ell_{k}}OH_{\ell_{1}}\cdots H_{\ell_{k^{\prime}}}$. A priori, the coefficients $\gamma_{i}=\beta_{j}\beta_{j^{\prime}}$ can be complex-valued, but the underlying Hermitian structure of the time-evolved observable (i.e. symmetry under taking adjoints) ensures that complex-valued coefficients must cancel out and the total sum ranges over $m_{\mathrm{tot}}\leq m^{2}$ terms with real-valued coefficients $\gamma_{i}\in\mathbb{R}$.
Estimating $\mathrm{tr}\!\left(O(t)\rho\right)$ is thus well-approximated by a real-valued, linear combination of the expectation values of Pauli operators $Q_{i}$, i.e.,
@@
For $K=O(\log(r/\epsilon))$ and $r=\lVert H\rVert t$, the number of Pauli expectation values in this sum is upper bounded by
$m_{\mathrm{tot}}$ $\leq\left(\sum_{k=0}^{K}L^{k}\right)^{2r}\!\!=\mathcal{O}(L^{2Kr})$ (7)
$=\mathrm{poly}\left(\left(\frac{\lVert H\rVert tL}{\epsilon}\right)^{\lVert H\rVert t}\right),$ (8)
where $\|H\|$ denotes the operator (or spectral) norm of the Hamiltonian $H$. Here, it is important to note that, strictly speaking, the dependence on time is (weakly) superexponential, as is the dependence on the system size, which indirectly enters via $\lVert H\rVert$.
The maximum Pauli weight over all $Q_{i}$s is given by
$w_{\mathrm{max}}=\min(\mathcal{O}\left(2ww_{O}\lVert H\rVert t\log(\lVert H\rVert t)),n\right),$ (9)
where we additionally need to take the maximum Pauli weight $w_{O}$ of the Pauli expansion of the observable $O$ into account.
Therefore, estimating time dynamics can effectively be reduced to a pure measurement task, albeit with potentially many different measurement configurations (single-qubit Pauli measurements). Since the number of expectation values of Pauli operators to be estimated grows exponentially in the simulation time, this tradeoff only pays off for short simulation times.
Comparing these insights with the original algorithm for simulating quantum dynamics with a truncated Taylor series [32], we find a similar exponential scaling, although in quantum circuit depth instead of sample complexity as we observe here.
@@
It is important to note that since the Taylor expansions rely on powers of the Hamiltonian and, thus, powers of sums of Pauli operators, not all will be pairwise different. This is due to their group structure and the fact that Pauli operators square to the identity. Consequently, the effective number of terms in this sum can be significantly smaller than the provided upper bound. However, an exponential scaling in the chosen cutoff $K$ and number of segments $r$ cannot be fully avoided.
For time-evolved observables $O(t)$ in particular, using the well-known fact that the time evolution in the Heisenberg picture for time-independent Hamiltonians and observables results in well-behaved nested commutators resulting from the Baker–Campbell–Hausdorff formula (see e.g. ), we find that instead of concatenating two truncated Taylor series of the time evolution operator as in Eq. (4), we can use
@@
which is reminiscent of cluster expansion techniques [41] using nested commutators to decide which terms to estimate in a classical algorithm on product states and stressing that we are leveraging the group structure of the Pauli operators. The advantage of such a reformulation is that we no longer have to calculate all combinations of Pauli operators to obtain the $Q_{i}$ in Eq. (5), many of which will cancel, but can use the underlying structure to calculate only non-vanishing terms. The connection to well-understood cluster expansion techniques provides a straightforward path to enhancing such techniques with quantum resources. Instead of performing such an expansion and estimating the individual Pauli expectation values classically for product states as in Ref. [41], we can measure them on arbitrary quantum states using existing expansions. Since nested commutators are a reformulation of the Taylor series of $O(t)$, the same error bounds as in Appendix B apply. Central to cluster expansion techniques and extensions thereof, thorough discussions can be found, e.g., in Refs. [41, 49]. Knowledge of the structure of the function of interest can thus help reduce the amount of classical post-processing.
Instead of concatenating two Taylor approximations of $U(t)$ for $t>0$ into an approximation of $O(t)$, as we have presented above, we can also directly Taylor expand the target function, as one would do for more generic functions $F(H,t)$. How such a direct expansion would present itself and differences in error and the number of terms are discussed in Appendix B.
@@ II.3 Error bounds
An approximation of $\mathrm{tr}\!\left(F(H,t)\rho\right)$ via a truncated Taylor series followed by an estimation of the resulting linear combination of expectation values of Pauli operators is accompanied by two types of errors: A systematic error due to the truncation of the Taylor series and a sampling error obtained during the estimation of the Pauli expectation values. Both of these error types can be rigorously bounded.
@@ II.3.1 Systematic errors
We can upper-bound the operator norm difference between the truncated Taylor series and target matrix function $F(H,t)$ by employing Taylor’s remainder theorem as showcased in Appendix A for a direct Taylor expansion of the particular matrix-valued function $F(H,t)=O(t)$. When stitching together multiple approximations as in Eq. (4) above, where we have constructed an approximation of $O(t)$ using a truncated Taylor series for $U(t)$, we can also combine the individual approximation errors to bound the total error when approximating the target function. As in Ref. [27, Lemma 1], given an approximation $\widetilde{U}$ of $U$ with $\lVert U-\widetilde{U}\rVert\leq\epsilon$, we can redefine $\widetilde{U}=U+U_{\epsilon}$ with $\lVert U_{\epsilon}\rVert\leq\epsilon$. We then find
@@ II.3.2 Sampling error
Using state-of-the-art methods, the linear combination of unitaries of Eq. (5) could be implemented deterministically using block encodings in the spirit of Refs. [48, 32] or in a randomized fashion using the frameworks of Refs. [35, 27]. However, since we are dealing with sums of Pauli operators, simply performing Pauli basis measurements as outlined in Eq. (6) suffices, and no further quantum circuitry is required.
@@
where we further normalize the pre-factors $\gamma_{j}$ such that we arrive at a probability distribution $\{p_{i},P_{i}\}$.
Since we are concatenating two Taylor approximations of the time evolution operator $U(t)=\mathrm{e}^{\mathrm{i}Ht}$, where the 1-norm of the terms of the truncated Taylor series is upper bounded by $\mathrm{e}^{\lambda t}$, we have
$\lVert\gamma\rVert_{1}=\mathrm{poly}\left(\mathrm{e}^{\lambda t}\right),$ (13)
directly influencing the number of required samples. A tighter bound can be obtained by truncating the Taylor expansion of $\mathrm{e}^{\lambda t}$ at the same order that the Taylor series of $\mathrm{e}^{\mathrm{i}Ht}$ is truncated.
Here, $\lambda$ will implicitly scale with the system size and thus the number of terms in the Hamiltonian, resulting in an exponential scaling with respect to the system size, which is likely to be a limiting factor in practice.
These insights are reminiscent of the runtime for cluster expansion methods [41, Theorem 6].
Sampling which Pauli expectation value $\mathrm{tr}\!\left(P_{i}\rho\right)$ to measure in each circuit evaluation from this distribution results in an unbiased estimator of $\widetilde{O}(t)$. Using Hoeffding’s inequality, we then find that
@@
samples suffice to estimate the linear combination of Pauli operators up to an error of $\epsilon$ with probability at least $1-\delta$. Note that just estimating $\mathrm{tr}\!\left(O\rho\right)$ for the same $\epsilon$ and $\delta$ would require $N\geq 2\ln(2/\delta)\epsilon^{-2}$ samples, meaning that this method increases the number of required samples by a factor of $\lVert\gamma\rVert_{1}^{2}$. The advantage of an importance sampling procedure is its independence on the number of terms $m_{\mathrm{tot}}$. Further, an importance sampling, first of the Taylor order and then of the terms in the Hamiltonian, can avoid the need for classical preprocessing to construct an explicit expansion.
This increase in sample complexity and the consequential limit to short times is precisely what the quantum algorithm for a randomized truncated Taylor series presented in Ref. [35] avoids with quantum resources to obtain a controllable overhead of $\exp(\lambda^{2}t^{2}/r)$.
The estimation suggested here can be improved using structural knowledge of the target function, e.g., by incorporating the nested commutator structure and quantum resources saved by immediately returning zero for all complex-valued $\gamma_{j}$ when considering $O(t)$. Alternatively, grouping strategies [19], graph optimization techniques [20], fully randomized measurements [10, 11], or other dedicated readout schemes could be used with their corresponding sampling error bounds.
@@
Especially quantum algorithms using randomized classical shadows [10, 11, 12, 13, 51] are of particular interest. Since these strategies by default aim to estimate a set of $m$ $n$-qubit Pauli operators $P_{i}$, they constitute the prototypical “measure now, compute functions of interest later” mindset. Thus, even if the main objective of a quantum algorithm differs from those discussed here, one still has the measurement data to reuse and can formulate a sampling error bound given the number of samples.
The performances of these methods are shown mostly heuristically with an analytical upper bound on the number of measurements given by Ref. [52, Theorem 3]. It essentially states that for an error $\epsilon\in(0,1)$ and failure probability $\delta$, it suffices to (sequentially) use
$N=\frac{2}{\epsilon^{2}(1-\epsilon)}3^{w_{\mathrm{max}}}\log\frac{3m_{\mathrm{tot}}}{\delta}$ (15)
copies of $\rho$ to approximate each $\mathrm{tr}\!\left(P_{i}\rho\right)$ up to (additive) error $\epsilon$, where $w_{\mathrm{max}}$ denotes the maximum Pauli weight and $m_{\mathrm{tot}}$ denotes the number of Pauli expectation values to be estimated.
For the presented algorithm, these results lead to the following conclusion: controlling the error of each $Q_{i}$ in Eq. (6) to $\epsilon/\lVert\gamma_{i}\rVert_{1}$ and thus bounding the overall sampling error of $\widetilde{O}(t)$ by $\epsilon$ requires
@@
copies of $\rho$. From Eq. (8), we know that the number of expectation values to be estimated $m_{\mathrm{tot}}$ scales in the worst case (slightly) doubly exponentially with the system size, meaning that the system size is a limiting factor not only due to its influence on $\lVert\gamma\rVert_{1}$.
While the actual performance could be better by orders of magnitude, these methods usually perform best for many low-weight or few global terms in contrast to the Taylor series, which outputs many terms with intermediate to large Pauli weight. Having obtained measurement data for another primary task, we can reuse the measurement data to secondarily estimate matrix functions $(H,t)\mapsto F(H,t)$ up to an explicitly determinable error.